A circuit with a 30-V battery connected to a 3.75-Ω resistor is used to melt a block of ice. How many seconds will this resistor take to melt 0.430 kg of ice at 0o C? The latent heat of fusion (LF) for water is 33.5 × 104 J/kg. Recall that Q = mLF
Answer #1

Here ,

Voltage , V = 30 V

resistance , R = 3.75 Ohm

let the time taken to melt the ice is t

Now ,

energy needed to melt ice = energy given by resistor

mass * latent heat = t * V^2/R

0.43 * 33.5 *10^4 = t * 30^2/3.75

solving for t

t = 600 s

the time taken to melt the ice is 600 s